# The ALMA Survey of Gas Evolution of PROtoplanetary Disks (AGE-PRO): I. Program Overview and Summary of First Results

### v1

@@ 3.3 The choice of non-JvM correction
Figure 4: Example of the 1.3 mm continuum and moment zero images of key science lines observed in the AGE-PRO program. The full gallery of line images are shown in Ruiz-Rodriguez et al. (2025, in press); Deng et al. (2025, in press); Agurto-Gangas et al. (2025, in press). The white contours indicate the 3 $\sigma$ level.
@@ 4 Analysis
Figure 3 shows a gallery of 1.3 mm continuum and 12CO (2-1) moment zero images of the 30 AGE-PRO sources. In Figure 4, we show examples of continuum and key science lines of three disks, one from each star-forming region. Here we briefly describe our general approach to measure line fluxes, radii, and the gas and dust disk masses. More in-depth descriptions of the analysis processes can be found in individual AGE-PRO papers (Ruiz-Rodriguez et al., 2025, in press; Deng et al., 2025, in press; Agurto-Gangas et al., 2025, in press; Trapman et al., in press, 2025a; Vioque et al., 2025, in press; Trapman et al., in press, 2025b).
@@ 4.2 Measuring line fluxes and radii
Figure 8: Comparison of disk properties in the Ophiuchus, Lupus, and Upper Sco samples. The circles are individual values and the squares are median values listed in Table 4. The gas disk sizes, RCO,90% are measured as radius that encloses 90 % of the 12CO (2-1). Due to the severe cloud/envelope contamination in the 12CO emission of the Ophiuchus disks, we only measure RCO,90% for the Lupus and Upper Sco disks (Trapman et al., in press, 2025b). The CO abundances of Ophiuchus disks are assumed to be at the ISM level (Zhang et al., 2020; Trapman et al., in press, 2025a) rather than constrained by N2H+, and therefore, are not plotted. The spectral index of Ophiuchus disks is not constrained due to the lack of Band 7 observations.
The gas disk sizes in Lupus and Upper Sco were measured as the radius that encloses 90 % of the 12CO flux in the integrated intensity (moment zero) map. As can be seen in Figure 3, 12CO emission is only marginally resolved for a large fraction of the Lupus and Upper Sco disks. As a result, gas disk sizes measured directly from the integrated intensity maps will be dominated by the size of the beam of the observations. To minimize this effect, AGE-PRO XI (Trapman et al., in press, 2025b) fit the 12CO integrated intensity maps with a 2D Nuker profile that was convolved with the clean beam of the observations, and measures gas disk sizes from the unconvolved best fit Nuker profiles using the curve-of-growth method described previously. We adopt the 90% flux CO sizes based on best-fitting model profile as the AGE-PRO recommended gas disk sizes (Trapman et al., in press, 2025b) and use these in all following discussions.
For the dust disk sizes, AGE-PRO X (Vioque et al., 2025, in press) employs visibility fitting to derive the surface brightness distribution of continuum emission and measure the 68%, 90%, and 95% flux radius for the 30 disks.
The final disk geometries established by AGE-PRO (inclinations and position angles) can be found in AGE-PRO X (Vioque et al., 2025, in press).
@@ 4.3 Gas and dust disk masses
The gas disk masses are estimated by comparing line fluxes with a large grid of thermo-chemical models. Please see AGE-PRO V (Trapman et al., in press, 2025a) for details. Here we only briefly outline the lines used for gas mass estimation. For Ophiuchus sources, the cloud contamination may affect the gas mass estimations, which is discussed in detail in AGE-PRO V (Trapman et al., in press, 2025a). We minimize the effect of cloud contamination by using the fluxes of the most optically thin C17O line measured within a tight Keplerian mask. For the majority of Ophiuchus disks, we only use C17O line fluxes for gas mass estimations. In cases of Oph 2, 6, and 9, their C18O emission also appears to be reasonably clear of cloud contamination, we compare C17O and C18O line fluxes with predictions of thermochemical models to constrain gas disk masses (Trapman et al., in press, 2025a). We assume the element C/H and O/H ratios are ISM level (correspondingly the ISM level of bulk CO gas abundance) in embedded disks, as suggested by previous study (Zhang et al., 2020).
For the Lupus and Upper Sco sources, we estimate the gas disk masses by comparing the 13CO, C18O, and N2H+ line fluxes, along with the 1.3 mm continuum fluxes, to a large grid of thermo-chemical models that compute the 2D temperature structure of the gas disk (Trapman et al., in press, 2025a). As discussed in §2.4.1, the N2H+-to-CO isotopologue flux ratios constrain the CO abundance ($x_{\rm CO}$), which, in turn, together with the line fluxes of CO isotopologues, provides constraints on the gas disk mass.
@@ 6 Overview of the AGE-PRO first results
The main goals of the AGE-PRO program are to measure the gas disk masses and sizes in three star-forming regions at different ages. In Figure 8, we show the individual and median disk properties in the three regions, including dust and gas disk masses and sizes, gas-to-dust mass and size ratios, CO abundances, and mm-continuum spectral indexes. The properties of individual disks are derived in AGE-PRO II-V and X-XI, and therefore these papers should be cited as the primary references for those properties (Ruiz-Rodriguez et al., 2025, in press; Deng et al., 2025, in press; Agurto-Gangas et al., 2025, in press; Trapman et al., in press, 2025a, in press, 2025b; Vioque et al., 2025, in press).
We calculate the median values for each region using the Kaplan–Meier estimator in the lifelines python package, to include upper and lower limits of individual data points. The median values, listed in Table 4, do not take into consideration survival bias, which is inherent in our selected sample as we are probing the disks that survived evolution and eventual dispersal. This bias is properly accounted for in population synthesis models of Tabone et al. (2025, in press).
@@ 6.1 Gas Evolution and Comparison with models
Gas disk mass: In each star-forming region, the inferred gas disk masses have a wide spread of $\sim$2 dex, as shown in Figure 8. The median gas mass of the Ophiuchus disk sample is one order of magnitude higher than that of the Lupus and Upper Sco regions. There is no significant difference between the masses of the Lupus and Upper Sco surviving disks. This suggests that most of the gas disk mass evolution happens at early times (between $<$1 Myr to 1-3 Myr). In our sample, there are seven, four, and five disks with gas masses higher than 1  MJup in the Ophiuchus, Lupus, and Upper Sco sub-samples, respectively. Please see individual gas masses and uncertainties in AGE-PRO III and V Deng et al. (2025, in press); Trapman et al. (in press, 2025a).
Gas disk size: The median gas disk size obtained from the 12CO increases from Lupus ($\sim$74 au) to Upper Sco ($\sim$110 au) (AGE-PRO XI Trapman et al., in press, 2025b). We note that Upper Sco disks are subject to mild external photoevaporation that may truncate gas disk sizes, as shown in (Anania et al., 2025, in press).
Comparison with population synthesis models: The gas disk masses and gas disk sizes obtained from AGE-PRO are compared to population synthesis models of turbulent-accretion models, including internal photoevaporation vs. MHD wind-driven accretion models (AGE-PRO VII Tabone et al., 2025, in press). The turbulence-driven models fail to simultaneously reproduce the median accretion rate and the disk mass. The MHD-driven accretion models reproduce better median disk properties. These MHD models require initially compact disks (5-10 au) and a magnetic field that declines with time.
@@
CO abundance: In the AGE-PRO Class II sample, CO abundances with respect to H2 are inferred from comparing N2H+ and CO isotopologue fluxes to grids of DALI thermochemical models (AGE-PRO V Trapman et al., in press, 2025a). We find the majority of our Class II disks have inferred CO abundances around 10-5, suggesting an order of magnitude CO depletion with respect to the ISM ratio.
@@ 6.2 Dust evolution and Comparison with models
Dust disk size: the dust median sizes of the surviving disks, measured from visibility fitting of the 1.3 mm continuum emission, are similar across the three regions (AGE-PRO X Vioque et al., 2025, in press). The apparent lack of size evolution differs from the findings of Hendler et al. (2020) which, using shallower surveys but a much larger sample of disks, reported more compact dust disks in Upper Sco compared to Lupus. However, our median dust disk sizes of Lupus and Upper Sco regions are consistent with their values within 3 $\sigma$ (AGE-PRO X Vioque et al., 2025, in press). Sensitive and larger sample observations are needed to improve the statistics on the evolution of dust disk size.
Dust Substructures: the continuum visibility fitting indicates that dust substructures are present in half of the AGE-PRO sample, with a tentative increase in the detection of large cavity-type structures in older disks (AGE-PRO X Vioque et al., 2025, in press). Substructures are detected in the young sources as well, suggesting these can form in older embedded disks. This is consistent with the ALMA Large Program eDISK, which found more substructures in their Class I disks compared to Class 0 ones (Ohashi et al., 2023). Similarly, the CAMPOS survey reported continuum substructures in 6 Class I/FS disks (Hsieh et al., 2024).
Comparison with dust evolution models: The decrease in mm-fluxes (or dust disk mass) with age can be explained when dust radial drift has been partially halted in pressure bumps (AGE-PRO VI Kurtovic et al., 2025, in press).
@@
I Zhang et al. (2025, in press) Program Overview and Summary of First Results
II Ruiz-Rodriguez et al. (2025, in press) Dust and Gas Disk Properties in the Ophiuchus Star-forming Region
III Deng et al. (2025, in press) Dust and Gas Disk Properties in the Lupus Star-forming Region
IV Agurto-Gangas et al. (2025, in press) Dust and Gas Disk Properties in the Upper Sco Star-forming Region
V Trapman et al. (in press, 2025a) Protoplanetary Gas Disk Masses
VI Kurtovic et al. (2025, in press) Comparison of Dust Evolution Models to AGE-PRO Observations
VII Tabone et al. (2025, in press) Testing Accretion Mechanisms from Disk Population Synthesis
VIII Anania et al. (2025, in press) Impact of External Photoevaporation on Disk Masses and Radii
@@
X Vioque et al. (2025, in press) Dust Substructures, Disk Geometries, and Dust Disk Radii
XI Trapman et al. (in press, 2025b) Beam-corrected Gas Disk Sizes for the AGE-PRO sample
XII Miley et al. (2025, in press) Extreme mm Variability in an Upper Sco Class II Disk
@@ 6.3.1 Disk properties in individual star-forming regions
AGE-PRO II (Ruiz-Rodriguez et al., 2025, in press), III (Deng et al., 2025, in press), and IV (Agurto-Gangas et al., 2025, in press) offer detailed observational results and discuss the statistical properties of AGE-PRO disks in the Ophiuchus, Lupus, and Upper Sco star-forming regions, respectively. These papers provide comprehensive analyses of the disks in each region, including statistical tests on the representativeness of the AGE-PRO sample, as well as detailed information on line emission morphology and flux distributions. Below, we highlight a key result from each paper.
Based on the morphology of 1.3 mm continuum images, AGE-PRO II (Ruiz-Rodriguez et al., 2025, in press) find 4 edge-on disks ($\geq$ 70 deg), and 3 highly inclined disks ($\geq$ 60 deg) in our Ophiuchus disk sample. For highly extincted objects, edge-on disks can potentially be misclassified as Class I/FS. Ruiz-Rodriguez et al. (2025, in press) provide new spectral type estimations and consider extinction for individual sources. The results confirm that 8 out of 10 objects are Class I/FS objects, while Oph 8 and 9 might belong to a more evolved stage, but not necessarily can be classified as Class II objects.
AGE-PRO III (Deng et al., 2025, in press) compares the gas masses of the Lupus disk sample estimated from the literature (Miotello et al., 2017), theoretical models by Ruaud et al. (2022), and gas masses from AGE-PRO (AGE-PRO V, Trapman et al., in press, 2025a). A good agreement in gas mass estimates, for large and massive disks, is found between the AGE-PRO estimates and those from the Ruaud et al. (2022) model grids, with both being significantly higher than the estimates from Miotello et al. (2017). For smaller disks, the two estimates show more variations but remain within the uncertainties.
AGE-PRO IV (Agurto-Gangas et al., 2025, in press) finds moderate correlations between the line fluxes of CO isotopologues and 1.3 mm continuum flux, suggesting a similar CO gas-to-dust mass ratio across the Upper Sco sample, with a larger scatter compared to the younger Lupus sample. The strong correlation between C18O and N2H+ fluxes is nearly identical for Lupus and Upper Sco, suggesting that $x_{\rm{CO}}$ is constant over the Class II stage of disk evolution. Further, the solid mass budget has diminished in older disks, while dust disk sizes and mm spectral indexes are, on average, comparable or even larger than that in the younger regions. This may be explained by the presence of dust traps impeding radial drift in these older disks.
@@ 6.3.2 Measurements of disk properties
AGE-PRO V (Trapman et al., in press, 2025a) describes how gas disk masses of the 30 AGE-PRO sources are constrained. A large grid of thermochemical models is employed to simultaneously fit the observed 1.3 mm continuum, CO isotopologue, and N2H+ line fluxes, as well as the 90% radii of 12CO line emission. The results showed that the median gas mass of the three regions appears to decrease by one order of magnitude from Ophiuchus to Lupus sample, but the median gas disk masses of the Lupus and Upper Sco samples are comparable.
AGE-PRO X (Vioque et al., 2025, in press) performs visibility fitting to the 1.3 mm continuum of these 30 disks. Substructures are identified in 15 out of the 30 disks, with 5 disks having large ($>$15 au) inner disk cavities. They discuss the evolution of dust disk radii and substructures among AGE-PRO disks across the three star-forming regions. A tentative trend is found, where the occurrence rate of disks with a large central cavity increases with age. In addition, Vioque et al. (2025, in press) reports the disk geometries (inclination angle and position angle) and dust disk sizes (Rdust,68% and Rdust,90%, Rdust,95%) used by AGE-PRO.
AGE-PRO XI (Trapman et al., in press, 2025b) measures the gas disk sizes for the 20 disks in our Lupus and Upper Sco samples, by fitting 12CO moment zero maps with beam-convolved models. The results show that median gas disk sizes are located at 74 au in Lupus and 110 au in Upper Sco. Combined with dust disk sizes measured in Vioque et al. (2025, in press), we also find the gas-to-dust size ratios are between 1-5.5. Contrary to models of dust evolution that predict an increasing size ratio with time (e.g., Birnstiel et al., 2010; Trapman et al., 2019), the AGE-PRO results show that the younger disks in Lupus have similar median ratios compared to the older disks in Upper Sco.

### v2

@@ 3.3 The choice of non-JvM correction
Figure 4: Example of the 1.3 mm continuum and moment zero images of key science lines observed in the AGE-PRO program. The full gallery of line images are shown in Ruiz-Rodriguez et al. (2025); Deng et al. (2025); Agurto-Gangas et al. (2025). The white contours indicate the 3 $\sigma$ level.
@@ 4 Analysis
Figure 3 shows a gallery of 1.3 mm continuum and 12CO (2-1) moment zero images of the 30 AGE-PRO sources. In Figure 4, we show examples of continuum and key science lines of three disks, one from each star-forming region. Here we briefly describe our general approach to measure line fluxes, radii, and the gas and dust disk masses. More in-depth descriptions of the analysis processes can be found in individual AGE-PRO papers (Ruiz-Rodriguez et al., 2025; Deng et al., 2025; Agurto-Gangas et al., 2025; Trapman et al., 2025a; Vioque et al., 2025; Trapman et al., 2025b).
@@ 4.2 Measuring line fluxes and radii
Figure 8: Comparison of disk properties in the Ophiuchus, Lupus, and Upper Sco samples. The circles are individual values and the squares are median values listed in Table 4. The gas disk sizes, RCO,90% are measured as radius that encloses 90 % of the 12CO (2-1). Due to the severe cloud/envelope contamination in the 12CO emission of the Ophiuchus disks, we only measure RCO,90% for the Lupus and Upper Sco disks (Trapman et al., 2025b). The CO abundances of Ophiuchus disks are assumed to be at the ISM level (Zhang et al., 2020; Trapman et al., 2025a) rather than constrained by N2H+, and therefore, are not plotted. The spectral index of Ophiuchus disks is not constrained due to the lack of Band 7 observations.
The gas disk sizes in Lupus and Upper Sco were measured as the radius that encloses 90 % of the 12CO flux in the integrated intensity (moment zero) map. As can be seen in Figure 3, 12CO emission is only marginally resolved for a large fraction of the Lupus and Upper Sco disks. As a result, gas disk sizes measured directly from the integrated intensity maps will be dominated by the size of the beam of the observations. To minimize this effect, AGE-PRO XI (Trapman et al., 2025b) fit the 12CO integrated intensity maps with a 2D Nuker profile that was convolved with the clean beam of the observations, and measures gas disk sizes from the unconvolved best fit Nuker profiles using the curve-of-growth method described previously. We adopt the 90% flux CO sizes based on best-fitting model profile as the AGE-PRO recommended gas disk sizes (Trapman et al., 2025b) and use these in all following discussions.
For the dust disk sizes, AGE-PRO X (Vioque et al., 2025) employs visibility fitting to derive the surface brightness distribution of continuum emission and measure the 68%, 90%, and 95% flux radius for the 30 disks.
The final disk geometries established by AGE-PRO (inclinations and position angles) can be found in AGE-PRO X (Vioque et al., 2025).
@@ 4.3 Gas and dust disk masses
The gas disk masses are estimated by comparing line fluxes with a large grid of thermo-chemical models. Please see AGE-PRO V (Trapman et al., 2025a) for details. Here we only briefly outline the lines used for gas mass estimation. For Ophiuchus sources, the cloud contamination may affect the gas mass estimations, which is discussed in detail in AGE-PRO V (Trapman et al., 2025a). We minimize the effect of cloud contamination by using the fluxes of the most optically thin C17O line measured within a tight Keplerian mask. For the majority of Ophiuchus disks, we only use C17O line fluxes for gas mass estimations. In cases of Oph 2, 6, and 9, their C18O emission also appears to be reasonably clear of cloud contamination, we compare C17O and C18O line fluxes with predictions of thermochemical models to constrain gas disk masses (Trapman et al., 2025a). We assume the element C/H and O/H ratios are ISM level (correspondingly the ISM level of bulk CO gas abundance) in embedded disks, as suggested by previous study (Zhang et al., 2020).
For the Lupus and Upper Sco sources, we estimate the gas disk masses by comparing the 13CO, C18O, and N2H+ line fluxes, along with the 1.3 mm continuum fluxes, to a large grid of thermo-chemical models that compute the 2D temperature structure of the gas disk (Trapman et al., 2025a). As discussed in §2.4.1, the N2H+-to-CO isotopologue flux ratios constrain the CO abundance ($x_{\rm CO}$), which, in turn, together with the line fluxes of CO isotopologues, provides constraints on the gas disk mass.
@@ 6 Overview of the AGE-PRO first results
The main goals of the AGE-PRO program are to measure the gas disk masses and sizes in three star-forming regions at different ages. In Figure 8, we show the individual and median disk properties in the three regions, including dust and gas disk masses and sizes, gas-to-dust mass and size ratios, CO abundances, and mm-continuum spectral indexes. The properties of individual disks are derived in AGE-PRO II-V and X-XI, and therefore these papers should be cited as the primary references for those properties (Ruiz-Rodriguez et al., 2025; Deng et al., 2025; Agurto-Gangas et al., 2025; Trapman et al., 2025a, b; Vioque et al., 2025).
We calculate the median values for each region using the Kaplan–Meier estimator in the lifelines python package, to include upper and lower limits of individual data points. The median values, listed in Table 4, do not take into consideration survival bias, which is inherent in our selected sample as we are probing the disks that survived evolution and eventual dispersal. This bias is properly accounted for in population synthesis models of Tabone et al. (2025).
@@ 6.1 Gas Evolution and Comparison with models
Gas disk mass: In each star-forming region, the inferred gas disk masses have a wide spread of $\sim$2 dex, as shown in Figure 8. The median gas mass of the Ophiuchus disk sample is one order of magnitude higher than that of the Lupus and Upper Sco regions. There is no significant difference between the masses of the Lupus and Upper Sco surviving disks. This suggests that most of the gas disk mass evolution happens at early times (between <<1 Myr to 1-3 Myr). In our sample, there are seven, four, and five disks with gas masses higher than 1  MJup in the Ophiuchus, Lupus, and Upper Sco sub-samples, respectively. Please see individual gas masses and uncertainties in AGE-PRO III and V Deng et al. (2025); Trapman et al. (2025a).
Gas disk size: The median gas disk size obtained from the 12CO increases from Lupus ($\sim$74 au) to Upper Sco ($\sim$110 au) (AGE-PRO XI Trapman et al., 2025b). We note that Upper Sco disks are subject to mild external photoevaporation that may truncate gas disk sizes, as shown in (Anania et al., 2025).
Comparison with population synthesis models: The gas disk masses and gas disk sizes obtained from AGE-PRO are compared to population synthesis models of turbulent-accretion models, including internal photoevaporation vs. MHD wind-driven accretion models (AGE-PRO VII Tabone et al., 2025). The turbulence-driven models fail to simultaneously reproduce the median accretion rate and the disk mass. The MHD-driven accretion models reproduce better median disk properties. These MHD models require initially compact disks (5-10 au) and a magnetic field that declines with time.
@@
CO abundance: In the AGE-PRO Class II sample, CO abundances with respect to H2 are inferred from comparing N2H+ and CO isotopologue fluxes to grids of DALI thermochemical models (AGE-PRO V Trapman et al., 2025a). We find the majority of our Class II disks have inferred CO abundances around 10-5, suggesting an order of magnitude CO depletion with respect to the ISM ratio.
@@ 6.2 Dust evolution and Comparison with models
Dust disk size: the dust median sizes of the surviving disks, measured from visibility fitting of the 1.3 mm continuum emission, are similar across the three regions (AGE-PRO X Vioque et al., 2025). The apparent lack of size evolution differs from the findings of Hendler et al. (2020) which, using shallower surveys but a much larger sample of disks, reported more compact dust disks in Upper Sco compared to Lupus. However, our median dust disk sizes of Lupus and Upper Sco regions are consistent with their values within 3 $\sigma$ (AGE-PRO X Vioque et al., 2025). Sensitive and larger sample observations are needed to improve the statistics on the evolution of dust disk size.
Dust Substructures: the continuum visibility fitting indicates that dust substructures are present in half of the AGE-PRO sample, with a tentative increase in the detection of large cavity-type structures in older disks (AGE-PRO X Vioque et al., 2025). Substructures are detected in the young sources as well, suggesting these can form in older embedded disks. This is consistent with the ALMA Large Program eDISK, which found more substructures in their Class I disks compared to Class 0 ones (Ohashi et al., 2023). Similarly, the CAMPOS survey reported continuum substructures in 6 Class I/FS disks (Hsieh et al., 2024).
Comparison with dust evolution models: The decrease in mm-fluxes (or dust disk mass) with age can be explained when dust radial drift has been partially halted in pressure bumps (AGE-PRO VI Kurtovic et al., 2025).
@@
I Zhang et al. (2025) Program Overview and Summary of First Results
II Ruiz-Rodriguez et al. (2025) Dust and Gas Disk Properties in the Ophiuchus Star-forming Region
III Deng et al. (2025) Dust and Gas Disk Properties in the Lupus Star-forming Region
IV Agurto-Gangas et al. (2025) Dust and Gas Disk Properties in the Upper Sco Star-forming Region
V Trapman et al. (2025a) Protoplanetary Gas Disk Masses
VI Kurtovic et al. (2025) Comparison of Dust Evolution Models to AGE-PRO Observations
VII Tabone et al. (2025) Testing Accretion Mechanisms from Disk Population Synthesis
VIII Anania et al. (2025) Impact of External Photoevaporation on Disk Masses and Radii
@@
X Vioque et al. (2025) Dust Substructures, Disk Geometries, and Dust Disk Radii
XI Trapman et al. (2025b) Beam-corrected Gas Disk Sizes for the AGE-PRO sample
XII Miley et al. (2025) Extreme mm Variability in an Upper Sco Class II Disk
@@ 6.3.1 Disk properties in individual star-forming regions
AGE-PRO II (Ruiz-Rodriguez et al., 2025), III (Deng et al., 2025), and IV (Agurto-Gangas et al., 2025) offer detailed observational results and discuss the statistical properties of AGE-PRO disks in the Ophiuchus, Lupus, and Upper Sco star-forming regions, respectively. These papers provide comprehensive analyses of the disks in each region, including statistical tests on the representativeness of the AGE-PRO sample, as well as detailed information on line emission morphology and flux distributions. Below, we highlight a key result from each paper.
Based on the morphology of 1.3 mm continuum images, AGE-PRO II (Ruiz-Rodriguez et al., 2025) find 4 edge-on disks ($\geq$ 70 deg), and 3 highly inclined disks ($\geq$ 60 deg) in our Ophiuchus disk sample. For highly extincted objects, edge-on disks can potentially be misclassified as Class I/FS. Ruiz-Rodriguez et al. (2025) provide new spectral type estimations and consider extinction for individual sources. The results confirm that 8 out of 10 objects are Class I/FS objects, while Oph 8 and 9 might belong to a more evolved stage, but not necessarily can be classified as Class II objects.
AGE-PRO III (Deng et al., 2025) compares the gas masses of the Lupus disk sample estimated from the literature (Miotello et al., 2017), theoretical models by Ruaud et al. (2022), and gas masses from AGE-PRO (AGE-PRO V, Trapman et al., 2025a). A good agreement in gas mass estimates, for large and massive disks, is found between the AGE-PRO estimates and those from the Ruaud et al. (2022) model grids, with both being significantly higher than the estimates from Miotello et al. (2017). For smaller disks, the two estimates show more variations but remain within the uncertainties.
AGE-PRO IV (Agurto-Gangas et al., 2025) finds moderate correlations between the line fluxes of CO isotopologues and 1.3 mm continuum flux, suggesting a similar CO gas-to-dust mass ratio across the Upper Sco sample, with a larger scatter compared to the younger Lupus sample. The strong correlation between C18O and N2H+ fluxes is nearly identical for Lupus and Upper Sco, suggesting that $x_{\rm{CO}}$ is constant over the Class II stage of disk evolution. Further, the solid mass budget has diminished in older disks, while dust disk sizes and mm spectral indexes are, on average, comparable or even larger than that in the younger regions. This may be explained by the presence of dust traps impeding radial drift in these older disks.
@@ 6.3.2 Measurements of disk properties
AGE-PRO V (Trapman et al., 2025a) describes how gas disk masses of the 30 AGE-PRO sources are constrained. A large grid of thermochemical models is employed to simultaneously fit the observed 1.3 mm continuum, CO isotopologue, and N2H+ line fluxes, as well as the 90% radii of 12CO line emission. The results showed that the median gas mass of the three regions appears to decrease by one order of magnitude from Ophiuchus to Lupus sample, but the median gas disk masses of the Lupus and Upper Sco samples are comparable.
AGE-PRO X (Vioque et al., 2025) performs visibility fitting to the 1.3 mm continuum of these 30 disks. Substructures are identified in 15 out of the 30 disks, with 5 disks having large (>>15 au) inner disk cavities. They discuss the evolution of dust disk radii and substructures among AGE-PRO disks across the three star-forming regions. A tentative trend is found, where the occurrence rate of disks with a large central cavity increases with age. In addition, Vioque et al. (2025) reports the disk geometries (inclination angle and position angle) and dust disk sizes (Rdust,68% and Rdust,90%, Rdust,95%) used by AGE-PRO.
AGE-PRO XI (Trapman et al., 2025b) measures the gas disk sizes for the 20 disks in our Lupus and Upper Sco samples, by fitting 12CO moment zero maps with beam-convolved models. The results show that median gas disk sizes are located at 74 au in Lupus and 110 au in Upper Sco. Combined with dust disk sizes measured in Vioque et al. (2025), we also find the gas-to-dust size ratios are between 1-5.5. Contrary to models of dust evolution that predict an increasing size ratio with time (e.g., Birnstiel et al., 2010; Trapman et al., 2019), the AGE-PRO results show that the younger disks in Lupus have similar median ratios compared to the older disks in Upper Sco.